Notson, Inc. produces several models of clocks. An outside supplier has offered to produce the commercial clocks for Notson for $420 each. Notson needs 1,200 clocks annually. Notson has provided the following unit costs for its commercial clocks:

Direct materials $100
Direct labor 140
Variable overhead 80
Fixed overhead (40% avoidable) 150

Prepare an incremental analysis which shows the effect of the make-or-buy decision.
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Make Buy Incremental
Direct Materials = 1200*100           120,000.00           120,000.00
Direct Materials = 1200*140           168,000.00           168,000.00
Variable Overhead = 1200*80             96,000.00             96,000.00
Relevant fixed overhead = 1200*150*40%             72,000.00             72,000.00
Purchase cost = 1200*420           504,000.00        (504,000.00)
Total relevant cost           456,000.00           504,000.00           (48,000.00)
Make is better option